OpenStudy (anonymous):

I will fan and metal!!! Which ordered pairs in the form (x, y) are solutions to the equation 7x – 5y = 28? Choose all answers that are correct. A. (−6, −14) B. (−1, −7) C. (4, 10) D. (7, 9)

OpenStudy (anonymous):

we can plug them in and see for exampole \((-4,-14)\) see if \[7\times (-6)+5\times (-14)=28\] yes or no?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

that one would be a YES http://www.wolframalpha.com/input/?i=7%28-6%29-5%28-14%29

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

lets try \((-1,-7)\) \[7\times (-1)-5\times (-7)=28?\] another YES http://www.wolframalpha.com/input/?i=7%28-1%29-5%28-7%29

OpenStudy (anonymous):

yw